OpenStudy (anonymous):

help pleaseeeeee!!! Bobby is bringing his collection of dimes and quarters to the bank to deposit into his savings account. If d represents the number of dimes in Bobby’s collection, q represents the number of quarters, and the collection is worth $63.75, which of the following equations models the value of Bobby’s collection? A. 63.75 = d + q B. 63.75 = 25d + 10q C. 63.75 = 10d + 25q D. 63.75 = 0.10d + 0.25q

OpenStudy (anonymous):

@ladiesman217

OpenStudy (anonymous):

I think that its D. 0.10d+0.25q because the others are more than 63.75 in total.

OpenStudy (anonymous):

@phi help

OpenStudy (phi):

First, try to understand the main idea. if you had 1 quarter how many dollars do you have ?

OpenStudy (anonymous):

75cent

OpenStudy (phi):

1 quarter is 25 cents

OpenStudy (phi):

can you say how many dollars 1 quarter is ?

OpenStudy (phi):

it will be a decimal

OpenStudy (anonymous):

so thw answer would be B

OpenStudy (phi):

1 quarter is $0.25

OpenStudy (phi):

if you have 2 quarters, how many dollars do you have ?

OpenStudy (anonymous):

oh so its D then

OpenStudy (phi):

if you have 2 quarters, how many dollars do you have ?

OpenStudy (anonymous):

0.50

OpenStudy (phi):

how did you figure that out ?

OpenStudy (anonymous):

lol i add it

OpenStudy (phi):

0.25+0.25 = 0.50 if you have 3 quarters it would be 0.25+0.25 +0.25 = 0.75 notice that instead of adding the same number 3 times, we can MULTIPLY 3* 0.25= 0.75 how much is 4 quarters worth (using multiplying)

OpenStudy (anonymous):

1.00 or 0.00390625

OpenStudy (phi):

just 1.00 if you added: 0.25+0.25++0.25+0.25= 1.00 if you multiply you get 4*0.25= 1.00

OpenStudy (phi):

If you had 8 quarters you would have 8*0.25 = 2.00 if you had 10 quarters you would have 10*0.25= 2.50

OpenStudy (anonymous):

63.75 = 0.10d + 0.25q

OpenStudy (phi):

now the new idea: if you have q quarters (q is some number that we don't know) how much is q quarters worth ?

OpenStudy (anonymous):

0.25

OpenStudy (phi):

I meant how much is q quarters worth

OpenStudy (anonymous):

1.00

OpenStudy (phi):

q might not be 4. You have to use the rule: multiply

OpenStudy (phi):

You write down q times 0.25 (we can't simplify to a single number because we don't know what q is)

OpenStudy (phi):

we follow the rule: 1 quarter = 1*0.25 2 quarters = 2*0.25 3 quarters = 3*0.25 q quarters = q*0.25 same idea with q as with 1, 2 or 3

OpenStudy (anonymous):

okay

OpenStudy (phi):

dimes are 10 cents which is 0.10 dollars if you have d dimes (d is some number), they are worth d*0.10 dollars if you have q quarters and d dimes, you have 0.10d + 0.25q dollars

OpenStudy (anonymous):

thanks